OpenStudy (anonymous):

Write an equation of the line that passes through points (3, 7) and (5, 11).

OpenStudy (australopithecus):

y = mx + b where m = slope b = y-intercept you can find the slope using these two points (x, y) (3, 7) (5, 11) To find the slope use the formula: \[ m = \frac{y_2- y_1}{x_2 - x_1}\] Then after you solve for m, sub m into the equation y = mx + b then insert one of your points into the equation and solve for b

OpenStudy (anonymous):

okay thank you :) but I don't think the possible answers are in y=mx+b format :/

OpenStudy (australopithecus):

All lines have the standard formula y = mx + b so follow my method and you will come to the right answer

OpenStudy (anonymous):

okay, thank you for your help :)